Week of 151123

WLCG Operations Call details

  • At CERN the meeting room is 513 R-068.

  • For remote participation we use the Vidyo system. Instructions can be found here.

General Information

  • The purpose of the meeting is:
    • to report significant operational issues (i.e. issues which can or did degrade experiment or site operations) which are ongoing or were resolved after the previous meeting;
    • to announce or schedule interventions at Tier-1 sites;
    • to inform about recent or upcoming changes in the experiment activities or systems having a visible impact on sites;
    • to provide important news about the middleware;
    • to communicate any other information considered interesting for WLCG operations.
  • The meeting should run from 15:00 until 15:20, exceptionally to 15:30.
  • The SCOD rota for the next few weeks is at ScodRota
  • General information about the WLCG Service can be accessed from the Operations Web
  • Whenever a particular topic needs to be discussed at the daily meeting requiring information from site or experiments, it is highly recommended to announce it by email to wlcg-operations@cernSPAMNOTNOSPAMPLEASE.ch to make sure that the relevant parties have the time to collect the required information or invite the right people at the meeting.

Links to Tier-1 downtimes

ALICE ATLAS CMS LHCB
  BNL FNAL  

Monday

Attendance:

  • local: Maria Alandes (SCOD), Maarten Litmaath (ALICE), Stefan Roiser (LHCb), Andrei Dumitru (DB)
  • remote: Jose Enrique Garcia Navarro (ATLAS), Christoph Wissing (CMS), Michael Ernst (BNL), Francesco Noferini (CNAF), Lisa Giachetti (FNAL), Rolf Rumler (IN2P3), Pavel (KIT), Christian (NDGF), Onno Zweers (NL-T1), Kyle Gross (OSG), Pepe Flix (pic), Tiju Idiculla (RAL)

Experiments round table:

  • CMS reports ( raw view) -
    • Very high load: 150k parallel jobs
    • No major issues

  • ALICE -
    • NTR

Sites / Services round table:

  • ASGC: NA
  • BNL: NTR
  • CNAF: We confirm the intervention on the general purpose router on the 24th (Tomorrow) at 18.00. We scheduled a down of about 2 hours (conservative approach because of the problem we got in the past).
  • FNAL: NTR
  • GridPP: NA
  • IN2P3: NTR
  • JINR: NA
  • KISTI: NA
  • KIT: Ongoing problems with disk buffer space that is not released automatically. This affects ATLAS.
  • NDGF: 2 minutes outage in the LHC OPN connection at Denmark due to an operator mistake.
  • NL-T1: NTR
  • NRC-KI: NA
  • OSG: NTR. Absent next Thursday due to Thanksgiving holiday.
  • PIC: dCache downtime scheduled tomorrow that will be unavailable from 9h-15h. Stefan asks whether the CEs are now draining, Pepe confirms the long queue is now draining.
  • RAL: NTR
  • TRIUMF: NA

  • CERN batch and grid services: NA
  • CERN storage services: NA
  • Databases: NTR
  • GGUS: NA
  • Grid Monitoring: NA
  • MW Officer: NA

AOB:

Thursday

Attendance:

  • local: Maria Alandes (SCOD), Maarten Litmaath (ALICE), Nacho Barrientos (Grid and Batch)
  • remote: Matteo (CNAF), Rolf Rumler (IN2P3), Thomas Hartmann (KIT), Christian (NDGF), Andrew Pickford (NL-T1), Tiju Idiculla (RAL), Stefan Roiser (LHCb)

Experiments round table:

  • ALICE -
    • So far the start of the HI run has been OK at least from the grid perspective.

  • LHCb reports ( raw view) -
    • Data Processing:
      • Data processing of pp data at T0/T1/T2 sites finishing
      • Monte Carlo mostly at T0/T1/T2/T2D, user analysis at T0/1/2D sites

Sites / Services round table:

  • ASGC: NA
  • BNL: NA
  • CNAF: Router intervention finished successfully.
  • FNAL: NA
  • GridPP: NA
  • IN2P3: NTR
  • JINR: NA
  • KISTI: NA
  • KIT: NTR
  • NDGF: dCache security upgrade finished successfully.
  • NL-T1: There is a downtime in SARA on 01.12.2015 to upgrade CREAM CE and dCache. More details in GOCDB.
  • NRC-KI: NA
  • OSG: NA
  • PIC: NA
  • RAL: NTR
  • TRIUMF: NA

Maarten asks whether there is any date to move the new version to production. Nacho explains that there are no plans at the moment, he will contact the service manager for more details.

  • CERN storage services: NA
  • Databases: NA
  • GGUS: NA
  • Grid Monitoring: NA
  • MW Officer: NA

AOB:

Edit | Attach | Watch | Print version | History: r12 < r11 < r10 < r9 < r8 | Backlinks | Raw View | WYSIWYG | More topic actions
Topic revision: r12 - 2015-11-26 - MaartenLitmaath
 
    • Cern Search Icon Cern Search
    • TWiki Search Icon TWiki Search
    • Google Search Icon Google Search

    LCG All webs login

This site is powered by the TWiki collaboration platform Powered by PerlCopyright & 2008-2019 by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding TWiki? Send feedback